# A Generic Machine Learning Framework for Radio Frequency Fingerprinting

### v1

@@ 5.1.3 Data from Drones: Counter-UAS
RF captures from four identical (i.e. same manufacturer, model and year) small hobbyist drones, Chubory F89, were collected with a COTS SDR in two different settings. These drones use the 2.4 GHz WiFi spectrum for the uplink (controller to drone) and downlink (drone to controller such as video feed) with frequency hopping within. The uplink and downlink transmissions had a central frequency of 2.470GHz, processed bandwidth is approximately 28 MHz to cover several of the subbands (namely, on the uplink transmission) and the SDR sample rate was set at 56MHz. The dataset contains approximately 40,000 emissions per drone. Our objective is to demonstrate that ML-based RFF enables the identification of each individual drone controller from raw I/Q data, in scenarios where the transceivers are the same make and model. This goes well beyond broader classification tasks that might aim to only determine the type of the drone (e.g. based on traditional signal classification) or drone versus non-drone RF emissions. In this context, SEI can facilitate threat detection and inform the patterns of life of drone operators in a C-UAS context. These capabilities can offer a substantial operational advantage in various scenarios, inclusive of cueing C-UAS effectors and resources planning.
5.2 Model Architectures
@@
This section documents model architectures which were used for experimentation. We used architectures that compose of convolution [26], Long Short Term Memory (LSTM) [18], transformer [55] and fully connected operations. The selection is described below:
@@
    Block Convolutional Neural Network (BCNN): multiple blocks of one-dimensional CNNs, batch normalization [20], leaky rectified linear unit [34] and max pooling layers. Five convolutional blocks are followed by two linear layers with an optional softmax layer;
@@
    Bidirectional LSTM (BILSTM): a LSTM-based architecture [16]. Here we use the PyTorch implementation with the following parameters: $\mbox{input\_size}=2,\mbox{hidden\_size}=512,\mbox{num\_layers}=2,\mbox{bidirectional}=\mbox{True}$, with a linear layer attached at the end of the model;
@@
    GLFormer: a transformer-based architecture [10] designed for SEI on Spaceborne AIS data. The architecture employs the following mechanisms: sequence embedding, Gated Attention Units (GAUs), Local Self-Attention (LSA), Sliding Local Self-Attention (SLA) [40] and Global Average Pooling (GAP) [31];
@@
For RFEC, we used four different auto-encoder architectures which we named: simpleAE, verysimpleAE, simpleconv1DAE and vanillaAE. The simpleAE architecture is a four layer fully connected network across both the encoder and decoder with leaky ReLU activation after each layer. The verysimpleAE model is the same as simpleAE but with only one layer, simpleconv1DAE parameterizes the encoder with three one-dimensional convolutional layers with batch normalization and leaky ReLU activation followed by three layers of fully connected networks in the encoder and decoder. The vanillaAE architecture is another fully connected network but with three layers in the encoder and decoder as well as batch normalization and leaky ReLU activation.
5.3 Setup
For all three datasets described in Section 5.1, and for all three tasks (SEI, EDA and RFEC), we use raw I/Q as our input representations. We also split each dataset into training, validation and testing datasets. In particular, $\mathcal{D}^{(t)}=\{\mathcal{D}^{(t)}_{train},\mathcal{D}^{(t)}_{valid},\mathcal{D}^{(t)}_{test}\}$ where $|\mathcal{D}^{(t)}_{train}|=1-p_{v}-p_{t}$, $|\mathcal{D}^{(t)}_{valid}|=p_{v}$ and $|\mathcal{D}^{(t)}_{test}|=p_{t}$. In our experiments, we choose $p_{v}=p_{t}=0.1$. Across all datasets, we use the multi-class cross-entropy loss function for SEI, the contrastive loss function for EDA and the mean-squared error loss for RFEC. For a single data sample, these are summarised as:
@@ 5.3 Setup
where $y_{j}$ is the ground-truth probability for emitter class $j$, $\hat{y}_{j}$ the predicted probability for the $j$th emitter class, $y$ and $\hat{y}$ in $\mathcal{L}^{(a)}$ are the ground-truth match and predicted match respectively, $\bf{y}$} and $^{\bf{\hat{y}}}$ are the ground-truth data and encoding of the data vectors in $\mathcal{L}^{(r)}$, $m$ is the margin parameter in the contrastive loss function, and $\lvert\lvert\cdot\rvert\rvert_{2}$ is the $L_{2}$ norm. In order to update the parameters which minimize these loss functions in practice, across the appropriate datasets, we define the iterative parameter update formula:
@@
where $\xi_{t}^{k}=(\phi_{t}^{k},\theta_{t}^{k})$, with $k$ referring to the current iteration, $t$ the task, $\tau^{(t)}_{k}$ the learning rate, ${\bf{\hat{Y}}}_{k}^{(t)}$ the model prediction, ${\bf{Y}}^{(t)}$ the ground truth label set and $\mathcal{L}^{(t)}$ the loss function. Here, we adopt mini-batch (MB) gradient descent to update the model parameters. Algorithm 1 shows pseudo-code for MB-Single Task Learning (MB-STL), which is executed for each individual task considered here.
@@
  end for
end for
OUTPUT: $v$th epoch for early stopping parameter set $\xi_{v,B}^{*}$
@@
In our experimentation, we adopt the Adam optimizer method [1] as $\mathcal{U}$, set the batch size $b_{size}=512,128,128$ for SEI, EDA, and RFEC respectively and assume an epoch $k$ is the completion of $\left[\lvert\mathcal{D}_{train}\rvert/b\right]$ batch updates, where $\left[\cdot\right]$ is the ceiling function. We choose the number of total epochs to be $E=200$ unless specified otherwise. Once the model parameter update is calculated via optimizing the loss function, it is mapped into a new parameter update that also minimizes the $L_{2}$ norm of the parameters $\xi$. In PyTorch, this is defined as a weight decay parameter and we set it to be $0.0005$. We choose the learning rate to be constant across training, selecting $\tau^{(i)}=7.5\times 10^{-3},1\times 10^{-3}$ (depending on model architecture), $\tau^{(a)}=1\times 10^{-3}$ and $\tau^{(r)}=1\times 10^{-3}$; all chosen empirically. In addition, we employ early stopping criteria using the validation dataset $\mathcal{D}_{valid}$, which retains the model with the best performance on the validation dataset across all epochs.
5.4 Results
In this section, experimental results found from solving CS-SEI, EDA, RFEC and OSR-EDA tasks on the datasets described in Section 5.1 are documented, using the experimental setup above. Across the experiments, where applicable, we display t-distributed stochastic neighbor embedding (t-SNE) plots [35] and confusion matrices, as well as quantitative metrics such as accuracy, precision, recall, F1 Score and Silhouette scores on K-Means clusters when clustering unsupervised data. The results section is structured by considering each task sequentially for different datasets. It is noted that the learnt RF fingerprints (referred to as embeddings) are typically of dimension 100 or above. However, for visualisation purposes we use t-SNE as a dimensionality reduction method in order to display the fingerprint in 2-D. Other dimensionality reduction methods such as UMAP [36] can be also used in lieu of t-SNE.
5.4.1 Use Case 1: CS-SEI for DMR Data
In this section, we document experimental work involving the DMR dataset for the CS-SEI task. Table 1 provides a summary of metrics (Accuracy, F1 Score, Precision and Recall) when evaluating trained BCNN, BILSTM, VGG19, GLFormer and FCN model architectures on the test dataset. Figures 3 and 4 display t-SNE plots and confusion matrices respectively on the test dataset. Different colours in the t-SNE plots refer to a different emitter class; notice how across all model architectures many of the emitter classes are separate with some overlap. The embeddings which are overlapped are actually the same manufacturer of DMR but different handsets, which tells us that there are crossover features for a specific model but they can still, to some degree, be separately identified by their RF fingerprints. The overlapping emitter embeddings behaviour from the t-SNE plot can also be observed in Figure 4, where confusion matrices are displayed for all the trained model architectures on this task. Emitters (2,3), (4,5) and (6,7) are unique devices but with the same make and manufacturer, and all architectures have some difficulty in fully separated their class boundaries whereas, for the most part, emitters 0 and 1 can be distinguished with better accuracy.
@@ 5.4.1 Use Case 1: CS-SEI for DMR Data
Table 1: Accuracy, F1 score, precision and recall of each trained model architecture evaluated on test dataset for SEI (8 handsets)
@@
5.4.2 Use Case 2: CS-SEI for Drone RF Data
In this section, we document experimental results involving the RF drone dataset for the CS-SEI task. Figure 5 shows a t-SNE embeddings plot of the drone test dataset and a confusion matrix, when evaluating it with a trained FCN CS-SEI model on the RF drone dataset. Notice how the embeddings are separate with some overlap and the confusion matrix shows consistently good performance across the drone classes.
Figure 5: Drone RF links test dataset evaluation. Top: t-SNE embeddings using FCN. Bottom: confusion matrix.
5.4.3 Use Case 3: EDA for DMR Data
In this section, we document experimental results involving the DMR dataset for the EDA task. Figure 6 shows four confusion matrices outlining trained model performance on the EDA test dataset for four separate architectures (VGG19, BCNN, GLFormer, BILSTM). Table 2 lists the performance metrics. The results show consistent good performance and similar performance across all model architectures.
@@ 5.4.3 Use Case 3: EDA for DMR Data
Table 2: Accuracy, F1 score, precision and recall of of trained BCNN, BILSTM, VGG19, GLFormer and FCN model architectures when evaluated on the EDA test dataset.
5.4.4 Use Case 4: EDA for Spaceborne AIS Data
Here, we present experimental results involving the AIS dataset collected by LEO satellites for the EDA task. Data from all of the 48 AIS emitters are used during the training and testing procedure, considering BCNN, BILSTM and FCN as model architectures. Figure 7 shows confusion matrices for three model architectures: BCNN, BILSTM and FCN. Similar to the other experiments, model performance is consistent across model architectures and demonstrates that ML-based RFF can deliver a high performance EDA, including for spaceborne maritime surveillance applications.
To assess the impact of noise on EDA performance, we synthetically add Additive White Gaussian Noise (AWGN) to the real RF captures from space, in order to generate datasets which have lower SNR. Figure 8 shows an accuracy plot of a BCNN-based EDA model against SNR. The typical trend can be seen: accuracy declines as SNR deteriorates. It is interesting to note that relying on decoding AIS burst typically requires SNRs $\gg 2dB$ whereas ML-based EDA can be accomplished at lower SNRs [61].
@@ 5.4.4 Use Case 4: EDA for Spaceborne AIS Data
Figure 8: Accuracy vs SNR for EDA task on spaceborne AIS data.
@@
Figure 9: t-SNE plot of training dataset embeddings of DMR dataset following Emitter Representation Learning. Top: simpleAE (L), simpleconv1DAE (R). Bottom: vanillaAE (L), verysimpleAE (R).
@@
Figure 10: K-Means Silhouette scores of training dataset embeddings. Top: simpleAE (L), simpleconv1DAE (R). Bottom: vanillaAE (L), verysimpleAE (R).
5.4.5 Use Case 5: RFEC for DMR Data
@@
This section presents experimental results for RFEC using the DMR dataset. Figure 9 shows t-SNE plots of the data embeddings for the training dataset using four model architectures. Notice how data structures have formed from learning lower dimensional representations of the emitter DMR via the autoencoder method. Figure 10 shows silhouette scores [46], which were computed on the training data embeddings, across a range of cluster numbers. It provides inconclusive evidence of emitters clusters being formed following unsupervised training. Nevertheless, it is included to demonstrate the RFEC task. It is possible that the DMR data has complexities which cannot be resolved via unsupervised approaches.
@@
Figure 11: Held out emitter data evaluation $p=0.5$: BCNN t-SNE embeddings (T), matched (red) and unmatched (blue) input pair distance distributions (M), K-Means Silhouette score for $K=2,3,..,9$ (B).
@@
Figure 12: Held out emitter data evaluation $p=0.25$: BCNN t-SNE embeddings (T), matched (red) and unmatched (blue) input pair distance distributions (M), K-Means Silhouette score for $K=2,3,..,9$ (B).
@@
Figure 13: Held out emitter data evaluation $p=0.025$: BCNN t-SNE embeddings (T), matched (red) and unmatched (blue) input pair distance distributions (M), K-Means Silhouette Score for $K=2,3,..,9$ (B).
5.4.6 Use Case 6: OSR-EDA for DMR data

### v2

@@ V-A3 RF Data from Drones: Counter-UAS
RF captures from four identical (i.e. same manufacturer, model and year) small hobbyist drones, Chubory F89, were collected with a COTS SDR in two different settings. These drones use the 2.4 GHz WiFi spectrum for the uplink (controller to drone) and downlink (drone to controller such as video feed) with frequency hopping within. The uplink and downlink transmissions had a central frequency of 2.47GHz, processed bandwidth is approximately 28 MHz to cover several of the subbands (namely, on the uplink transmission) and the SDR sampling rate was set at 56MHz. The dataset contains approximately 40,000 emissions per drone. Our objective is to demonstrate that ML-based RFF enables the identification of each individual drone controller from raw I/Q data, in scenarios where the transceivers are the same make and model. This goes well beyond broader classification tasks that might aim to only determine the type of the drone (e.g. based on traditional signal or modulation-scheme classification) or drone versus non-drone RF emissions. In this context, SEI can facilitate threat detection and pattern-of-life analysis of drone(s) operators in a C-UAS context. These capabilities can offer a substantial operational advantage in various scenarios, inclusive of cueing C-UAS effectors and resource planning.
V-B Model Architectures
@@
The experimentation conducted in this paper uses example common ML model architectures that compose of convolution [70], long short term memory (LSTM) [71], transformer [37] and fully connected operations. They are:
@@
    Block convolutional neural network (BCNN): multiple blocks of one-dimensional CNNs, batch normalisation [72], leaky rectified linear unit [73] and max pooling layers. Five convolutional blocks are followed by two linear layers with an optional softmax layer;
@@
    Bidirectional LSTM (BILSTM): a LSTM-based architecture [74]. Specifically, we use the PyTorch implementation with the following parameters: $\mbox{input\_size}=2,\mbox{hidden\_size}=512,\mbox{num\_layers}=2,\mbox{bidirectional}=\mbox{True}$, with a linear layer attached at the end of the model;
@@
    GLFormer: a transformer-based architecture for CS-SEI introduced in [22]. The architecture employs the following mechanisms: sequence embedding, gated attention units (GAUs), local self-attention (LSA), sliding local self-attention (SLA) [75] and global average pooling (GAP) [76];
@@
For RFEC, we used four different auto-encoder architectures which we named: simpleAE, verysimpleAE, simpleconv1DAE and vanillaAE. The former model is comprised of a four layer fully connected network across both the encoder and decoder with leaky ReLU activation after each layer. The verysimpleAE model is the same as simpleAE, however with only one layer. The simpleconv1DAE parameterises the encoder with three one-dimensional convolutional layers with batch normalisation and leaky ReLU activation followed by three layers of fully connected networks in the encoder and decoder. The vanillaAE architecture is another FCN, but with three layers in the encoder and decoder, batch normalisation and leaky ReLU activation. The standard K-means clustering algorithm is applied to the learnt low-dimensional representations (i.e. outputs of the encoder).
It is emphasised that other suitable ML model architectures can be applied within the presented generic RFF framework for any of the considered tasks.
V-C Setup
For all three datasets described in Section 5.1, and for all three tasks (SEI, EDA and RFEC), the input to the ML models is the received transmission raw I/Q data. We split each dataset into training, validation and testing data. In particular, $\mathcal{D}^{(t)}=\{\mathcal{D}^{(t)}_{train},\mathcal{D}^{(t)}_{valid},\mathcal{D}^{(t)}_{test}\}$ where $|\mathcal{D}^{(t)}_{train}|=1-p_{v}-p_{t}$, $|\mathcal{D}^{(t)}_{valid}|=p_{v}$ and $|\mathcal{D}^{(t)}_{test}|=p_{t}$ such we choose here $p_{v}=p_{t}=0.1$. Across all datasets, we use the multi-class cross-entropy loss function for SEI, the contrastive loss function for EDA and the mean-squared error loss for RFEC. For a single data sample, these are summarised as:
@@ V-C Setup
where $y_{j}$ is the ground-truth probability for emitter class $j$, $\hat{y}_{j}$ the predicted probability for the $j$th emitter class, $y$ and $\hat{y}$ in $\mathcal{L}^{(a)}$ are the ground-truth match and predicted match respectively, $\bf{y}$} and $^{\bf{\hat{y}}}$ are the ground-truth data and encoding of the data vectors in $\mathcal{L}^{(r)}$, $m$ is the margin parameter in the contrastive loss function, and $\lvert\lvert\cdot\rvert\rvert_{2}$ is the $L_{2}$ norm. In order to update the parameters which minimise these loss functions in practice, across the appropriate datasets, we define the iterative parameter update formula:
@@
where $\xi_{t}^{k}=(\phi_{t}^{k},\theta_{t}^{k})$, with $k$ referring to the current iteration, $t$ the task, $\tau^{(t)}_{k}$ the learning rate, ${\bf{\hat{Y}}}_{k}^{(t)}$ the model prediction, ${\bf{Y}}^{(t)}$ the ground-truth label set and $\mathcal{L}^{(t)}$ the loss function. Here, we adopt mini-batch (MB) gradient descent to update the model parameters. Algorithm 1 shows pseudo-code for MB-single task learning (MB-STL), which is executed for each individual task.
@@
OUTPUT: $\xi_{v,B}^{*}$ with an early stopping criterion
@@
We adopt the Adam optimiser method [78] as $\mathcal{U}$, set the batch size $b_{size}=512,128,128$ for SEI, EDA, and RFEC, respectively, and assume an epoch $k$ is the completion of $\lceil\lvert\mathcal{D}_{train}\rvert/b\rceil$ batch updates, where $\left\lceil\cdot\right\rceil$ is the ceiling function. We choose the number of total epochs to be $E=200$ unless specified otherwise. Once the model parameter update is calculated via optimising the loss function, it is mapped into a new parameter update that minimises the $L_{2}$ norm of the parameters $\xi$. This is defined as a weight decay parameter and it is set to $5\times 10^{-4}$. The learning rate is constant across training, with $\tau^{(i)}=7.5\times 10^{-3},1\times 10^{-3}$ (depending on model architecture), $\tau^{(a)}=1\times 10^{-3}$ and $\tau^{(r)}=1\times 10^{-3}$; all chosen empirically. We apply early stopping criteria using the validation dataset $\mathcal{D}_{valid}$, which retains the model with the best performance across all epochs. It is noted that other set ups and configurations, inclusive of loss function and optimiser, can be employed within the introduced ML framework.
@@
TABLE I: Accuracy, F1 score, precision and recall of each trained model architecture evaluated on test dataset for SEI (8 handsets)
@@
VI Results
Results for CS-SEI, EDA, RFEC and OSR-EDA from the datasets in Section V-A are presented here. Where applicable, we display confusion matrices and quantitative metrics such as accuracy, precision, recall, F1 Score and Silhouette scores on K-Means clusters. It is noted that the learnt RF fingerprints representations (i.e. embeddings) are typically of dimension 100 or above. For visualisation purposes, we use t-distributed stochastic neighbor embedding (t-SNE) [79] for dimensionality reduction to display the fingerprint in 2-D; others such as UMAP [80] can be employed.
VI-A Use Case 1: CS-SEI for DMR Data
Table I provides a summary of CS-SEI results (i.e. metrics accuracy, F1 score, precision and recall) with BCNN, BILSTM, VGG19, GLFormer and FCN model architectures for the DMR test dataset. The corresponding confusion matrices are depicted in Figure 3. It can be noticed that all of the examined ML models deliver comparable SEI performance, albeit the largest model VGG19 (i.e. in terms of model size and number of parameters) has a marginally higher accuracy compared to a substantially simpler-smaller architectures (e.g. FCN and BCNN). This concurs with results in [81].
Figure 4 displays the t-SNE plots (i.e. RF fingerprint representation in 2-D). Different colours in each plot refer to a different emitter class and each point within each plot is obtained from a data sample (i.e. time-series of I/Q RF signal from a given transmission for example of length 512). It can be noted, the RF fingerprint representations from the eight emitter classes are separable across all of the assessed ML models, despite some overlap. The most visible overlaps in fact belong to DMR handsets that are of the same model and make (i.e. from the same manufacturer). Whilst this is expected and implies that there are crossover features for a specific model, their emissions can still be individually identified by their RF fingerprints, especially if this is over several data samples. The overlapping emitter embeddings behaviour from the t-SNE plot can also be observed in Figure 3, where confusion matrices are displayed for all the trained model architectures on this task. Emitters (2,3), (4,5) and (6,7) are unique devices but with the same make and manufacturer, and all architectures have some difficulty in fully separating their class boundaries whereas, for the most part, emitters 0 and 1 can be distinguished with better accuracy.
VI-B Use Case 2: CS-SEI for Drone RF Data
Figure 5 shows the t-SNE embeddings plot and a confusion matrix of the trained FCN CS-SEI model for the drone RF links dataset in Section V-A3. The fingerprints for each of the four drones (i.e. same model and make, thus four different instances of the same transmitter ) are visibly separable with some overlap. The confusion matrix demonstrates consistently good performance across all classes, achieving an overall accuracy exceeding $96\%$ with the basic FLC model.
Figure 5: CS-SEI for the drone RF links test dataset. Top: t-SNE embeddings using FCN. Bottom: confusion matrix.
VI-C Use Case 3: EDA for DMR Data
The EDA results for the DMR dataset is depicted in Figure 6, displaying the four confusion matrices for the VGG19, BCNN, GLFormer and BILSTM models. Table II lists the performance metrics. Both shows good performance across all of the considered architectures.
@@ VI-C Use Case 3: EDA for DMR Data
TABLE II: EDA performance of trained BCNN, BILSTM, VGG19, GLFormer and FCN models for EDA dataset.
VI-D Use Case 4: EDA for Spaceborne AIS Data
Data from all of the 48 AIS emitters collected by the LEO satellites are used during the training and testing procedure of the three selected models: BCNN, BILSTM and FCN. Figure 7 displays their confusion matrices. Similar to the other experiments above, this figure illustrates that ML-based RFF can deliver a high accuracy EDA for various model architectures, including for spaceborne maritime surveillance.
To assess the impact of noise on the EDA performance, we synthetically add additive white Gaussian noise (AWGN) to the real RF captures from space. The outcome, accuracy versus SNR, is shown in Figure 8 for the BCNN-based EDA model. The typical trend can be seen, where accuracy declines as SNR decreases. It is interesting to note that relying on decoding AIS burst (i.e. associate AIS bursts based on the extracted MMSI) typically requires SNRs $\gg 2$ dB whereas ML-based EDA can be accomplished at significantly lower SNRs [82].
@@ VI-D Use Case 4: EDA for Spaceborne AIS Data
Figure 8: Accuracy versus SNR for EDA on spaceborne AIS data.
@@
Figure 9: t-SNE plot of DMR dataset embeddings following emitter representation learning step. Top: simpleAE (L), simpleconv1DAE (R). Bottom: vanillaAE (L), verysimpleAE (R).
@@
Figure 10: K-Means Silhouette scores of DMR dataset embeddings. Top: simpleAE (L), simpleconv1DAE (R). Bottom: vanillaAE (L), verysimpleAE (R).
VI-E Use Case 5: RFEC for DMR Data
@@
Figure 9 depicts the t-SNE plots of the embeddings (RF fingerprint representation) for four model architectures using the DMR dataset. Notice how data structures have formed from learning lower dimensional representations of the emitter with autoencoders. Figure 10 shows silhouette scores [83], which were computed on the training data embeddings, across a range of cluster numbers. It provides inconclusive evidence of emitters clusters being formed following unsupervised training. Nevertheless, it is included to demonstrate the RFEC task. It is possible that the DMR data has complexities which cannot be resolved via unsupervised approaches.
@@
Figure 11: DMR OS-EDA results $p=0.5$. Top: BCNN t-SNE embeddings, matched (red) and unmatched (blue) input. Middle: pair distance distributions. Bottom: K-Means Silhouette score.
@@
Figure 12: DMR OS-EDA results $p=0.25$. Top: BCNN t-SNE embeddings, matched (red) and unmatched (blue) input. Middle: pair distance distributions. Bottom: K-Means Silhouette score.
@@
Figure 13: DMR OS-EDA results $p=0.025$. Top: BCNN t-SNE embeddings, matched (red) and unmatched (blue) input. Middle: pair distance distributions. Bottom: K-Means Silhouette score.
VI-F Use Case 6: OS-EDA for DMR data